Cooper Movie Studio Corp. makes movies and is interested in lowering its operating costs for the following year, while maintaining the high quality and appeal of its movies. Cooper’s management is concerned about the additional costs the company would have to incur if new industry regulation is passed by Congress. The chart at the top of the next page contains a description of this top risk, an inherent risk assessment, three risk response alternatives, and a residual risk assessment for each response alternative.

Inherent Risk Risk Response Residual Risk
Risk Likelihood Impact
(on operating costs)
Alternatives Likelihood Impact
(on operating costs)
Poor behavior by others in the industry results in Congress passing costly new legislation that regulates the behavior of all movie studios.   30% $80,000,000 A—Share Cooper’s effective marketing processes with competitors to help mitigate poor industry behavior 20% $80,000,000
B—Lobby Congress on behalf of the movie industry generally and Cooper in particular 15% $50,000,000
C—Take no action in response to possible new regulation 30% $80,000,000

Finally, Cooper’s management accountants estimate that the incremental cost of implementing risk response A is $3,000,000 and the incremental cost of implementing risk response B is $13,000,000.

Required:

If an amount is zero, leave it blank or enter "0".

1. Calculate the inherent risk for Cooper.

$24000000

2. Calculate the residual risk for Cooper associated with each of the three risk response alternatives A, B, and C.

A. $16000000
B. $7500000
C. $24000000

3. Calculate the benefit for Cooper associated with each of the three risk response alternatives A, B, and C.

A. $8000000
B. $16500000
C. $0

4. Calculate the net benefit for Cooper associated with each of the three risk response alternatives A, B, and C.

A. $
B. $3500000
C. $0

5. Using net benefit as the criterion, which risk response should Cooper choose to implement?

The risk response alternative with the greatest net benefit is alternative A
Answer #1

1.Inherent risk for cooper = impact x likelihood =80000000 x0.3 = 24000000

2. Residual risk

A -(80000000x0.2) = 16000000

B - (50000000x0.15) = 75000000

C - (80000000x0.3) = 24000000

3. Benefit for Cooper

A - (-16000000)=8000000

B- (-7500000)=16500000

C- (-24000000)=0

4. Net Benefit for Cooper

A- 24000000-16000000 = 8000000

B- 24000000-7500000 =1650000

C- 24000000-24000000 =0

5. Option B is the most suitable, as risk is least